### v3

# Boost 3D Reconstruction using Diffusion-based Monocular Camera Calibration

Abstract
In this paper, we present DM-Calib, a diffusion-based approach for estimating pinhole camera intrinsic parameters from a single input image. Monocular camera calibration is essential for many 3D vision tasks. However, most existing methods depend on handcrafted assumptions or are constrained by limited training data, resulting in poor generalization across diverse real-world images. Recent advancements in stable diffusion models, trained on massive data, have shown the ability to generate high-quality images with varied characteristics. Emerging evidence indicates that these models implicitly capture the relationship between camera focal length and image content. Building on this insight, we explore how to leverage the powerful priors of diffusion models for monocular pinhole camera calibration. Specifically, we introduce a new image-based representation, termed Camera Image, which losslessly encodes the numerical camera intrinsics and integrates seamlessly with the diffusion framework. Using this representation, we reformulate the problem of estimating camera intrinsics as the generation of a dense Camera Image conditioned on an input image. By fine-tuning a stable diffusion model to generate a Camera Image from a single RGB input, we can extract camera intrinsics via a RANSAC operation. We further demonstrate that our monocular calibration method enhances performance across various 3D tasks, including zero-shot metric depth estimation, 3D metrology, pose estimation and sparse-view reconstruction. Extensive experiments on multiple public datasets show that our approach significantly outperforms baselines and provides broad benefits to 3D vision tasks.
1 Introduction
Camera calibration is a foundational task in 3D computer vision, critical for numerous applications such as camera pose estimation [58], 3D reconstruction [62], and zero-shot metric depth estimation [85]. Traditional methods primarily focus on multi-view calibration, typically involving multiple images of fixed intrinsics [47] or multiple images of checkerboard patterns [95].
However, these methods depend heavily on dense multi-view images with sufficient overlap, making them cumbersome and often impractical for sparse-view or even monocular setups. Consequently, monocular camera calibration has garnered significant research interest.
Monocular camera calibration is inherently an ill-posed problem, requiring additional information to address it. Traditional approaches have attempted to incorporate handcrafted knowledge to solve the problem, such as the gravity direction [71], Manhattan World constraints [39], and human face priors [25]. However, these handcrafted insights often fail to generalize effectively across large diverse real-world scenarios. To overcome these limitations, recent studies [96] have recast monocular camera calibration as a learning-based regression problem, leveraging a single image to directly infer its intrinsic parameters.
While learning-based methods benefit from data-driven knowledge, outperforming traditional approaches, they are constrained by the limited availability of public datasets. As a result, these methods tend to overfit on training data and exhibit poor generalization to unseen scenarios. This limitation raises a critical question: What kind of knowledge is necessary to develop a robust camera calibration method that exhibits strong generalization capabilities?
One promising solution lies in leveraging stable diffusion priors [53]. The intuition behind stems from a key observation: stable diffusion models possess an implicit understanding of imaging across different focal lengths. As is widely known, Cameras with long focal lengths tend to compress spatial relationships, resulting in a more flattened image perspective, while wide-angle cameras exaggerate depth and distance, resulting in a more pronounced perspective effect.
Figure 1: Images generated using text prompts that specify different focal lengths.
As illustrated in Fig. 1, we present two portrait images generated by a stable diffusion model [53] using similar text prompts but with varying focal length descriptions. Clearly, the left image prompted with “long focal length”, exhibits a more blurred background and shallower depth of field compared to the right image. This demonstrates that, by training on large-scale image-text pairs, these models encapsulate knowledge related to imaging characteristics associated with different focal lengths.
Despite these advancements, a key challenge persists: how to effectively leverage diffusion priors for high-precision camera calibration? In this paper, we set out to explore this question and introduce DM-Calib, a diffusion-based model for estimating intrinsic camera parameters from a single image. We recognize that the format used to encode camera intrinsics is essential for effective monocular camera calibration using diffusion models. To address this, we conduct in-depth investigation for various camera representations and develop the Camera Image, a novel image-based representation specifically engineered for seamlessly integration with pre-trained diffusion models, thereby preventing loss of information. Subsequently, we train a diffusion model that takes a single image as input and generates the Camera Image, followed by a RANSAC algorithm to solve the camera intrinsic parameters. Moreover, we demonstrate how to integrate the proposed camera calibration with diffusion-based metric depth estimation, which allows the recovery of true-scale depth measurements from a single image. Furthermore, our experiments show that the recovered camera calibration results significantly improve the performance of various downstream tasks, including camera pose estimation, sparse-view 3D reconstruction, and novel view synthesis, showcasing the robustness and effectiveness of DM-Calib in advancing accurate monouclar camera calibration.
To summarize, our main contributions are:
  • We introduce the Camera Image, a novel image-based representation specifically designed to encode camera intrinsic, optimized to use with pretrained diffusion models.
  • We present DM-Calib, a generative foundation model that provides highly accurate estimations of camera intrinsics. Additionally, it can seamlessly integrate with various downstream tasks, showcasing its effectiveness and robustness to images from various scenarios.
  • Extensive experiments on multiple public datasets show that our approach significantly outperforms baselines and provides broad benefits to 3D vision tasks.
2 Related Work
2.1 Monocular Camera Calibration
Calibrating camera intrinsics, also known as self-calibration, is one of the most fundamental problems in geometric computer vision. Geometric methods typically assume multiple input images with the same intrinsic matrix. These methods can be broadly classified as direct methods and stratified methods. Direct methods [88, 22, 43] solve the Kruppa’s equation [17] to obtain camera intrinsic parameters, which are generally more fragile to noises. In comparison, stratified methods [21, 69, 47] estimate camera intrinsics from a projective reconstruction by gradually recovering the affine and Euclidean structures such as the plane at infinity and absolute quadrics.
Traditional monocular camera calibration methods typically assume specific geometric structures to estimate intrinsics. For example, with Manhattan World assumption [12], camera intrinsics can be inferred from vanishing points [36, 57, 75], or by jointly estimating the horizon line [89, 63]. Other approaches rely on calibration objects such as checkerboards [95], line segments [94], spheres [91], pyramid frustums [29], or even human faces [25, 39]. Similarly,  Schops et al. [60], Grossberg and Nayar [20] parametrize intrinsics via per-pixel ray directions and solving that via capturing images of a specific calibration pattern. Despite satisfactory results, these methods are constrained by their reliance on specific objects, which limits their applicability in unconstrained, real-world scenarios. Recently, Zhu et al. [96] introduced incident maps to regress camera intrinsics, enabling the detection of geometric manipulations like cropping.
Apart from the above geometry-based methods, some works attempt to leverage the strong generative models for camera calibration. To the best of our knowledge, the approach by He et al. [23] is the only existing method that formulates camera intrinsic estimation as a generative task, leveraging incident maps with diffusion models. Although this method outperforms non-diffusion-based methods such as Zhu et al. [96] in unconstrained settings, we argue that the full potential of diffusion models remains under-utilized and that its performance hinges on joint training with depth images. In contrast, our approach introduces a camera representation that is more inherently compatible with diffusion models, thereby eliminating the need to generate non-textured incident maps or rely on the joint training of additional geometric information.
2.2 Diffusion Models in 3D tasks
Recently, diffusion models [24, 64] have emerged as powerful tools across various domains, particularly in the field of computer vision, where Text-to-Image diffusion models [55] and their extensions [48, 53, 54, 93, 28] have garnered significant attention. Compared to GAN-based approaches [5], several studies have highlighted the advantages of diffusion models, especially when used as prior geometric cues in 3D tasks. Notable examples include view synthesis [37, 40], camera calibration [23], normal estimation [81, 38], and depth estimation [15, 33].  Zhang et al. [92] modeled the camera as rays and a diffusion pipeline is used to determine its final direction. Although the method shows good results, it relies on multi-view data and is mainly suited for object-centric scenes, which limits its practical application. In this work, we focus on leveraging diffusion models for in-the-wild monocular camera intrinsic estimation, which serves as a foundation for enhancing a series of downstream tasks, such as monocular metric depth estimation, pose estimation, and sparse-view reconstruction.
2.3 Monocular Depth Estimation
For depth estimation, several works [33, 15, 27, 66] have shown that diffusion models can be fine-tuned to predict affine-invariant depth, achieving not only finer detail but also more accurate estimates than traditional methods [49, 83, 80, 84]. Recent studies [79, 81, 18] have highlighted that diffusion models can serve as pre-trained networks for deterministic, one-step inference. However, none of the current approaches leverage pre-trained diffusion models for metric depth prediction. Existing zero-shot monocular metric depth estimation methods, such as [4, 85, 46, 26], have demonstrated accurate results, yet they still face challenges in capturing geometric details and foreground-background relationships, particularly in outdoor environments. Moreover, these methods usually employ contrastive learning pretrained encoders (i.e., DINO [9]) or classification pretrained encoders [14], and randomly initialize the decoder, which are trained on fewer images compared to diffusion models. Building upon this, we extend diffusion models to metric depth estimation. To the best of our knowledge, our approach is the first to employ pre-trained diffusion models for metric depth estimation, achieving finer geometric detail and competitive performance across diverse benchmarks.
3 Method
Given a single input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times 3}$, our objective is to recover its camera intrinsic matrix $\bm{K}$}. To efficiently and losslessly integrate camera intrinsics prediction with diffusion models [53], we introduce Camera Image (Fig. 2) to encode camera intrinsics as a detail-preserving color image (see Sec. 3.2). We reformulate camera calibration as a conditional generation task, transforming the text-to-image (T2I) diffusion model into an image-to-Camera-Image (I2C) model, from which camera intrinsics are recovered via a RANSAC algorithm (see Sec. 3.3). Our proposed calibration method significantly boosts the performance of downstream 3D vision tasks, including monocular metric depth estimation, 3D reconstruction, and pose estimation (see Sec. 3.4). Before introducing our method, we first revisit the preliminary concepts related to diffusion models.
3.1 Preliminaries on Diffusion Model
Diffusion models (DMs) [24] learn to model a data distribution $p_{\text{data}}({\mathbf{x}})$ by progressively denoising a noise variable that is initially sampled from a normal distribution. Recognizing the efficiency issue associated with generating high-resolution images, Rombach et al. [53] introduced latent diffusion models (LDMs), which operate the diffusion process in the latent space of a pretrained variational autoencoder (VAE) [34] with an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$}.
[…]
Figure 2: Visualization of incidence map and Camera Image. We show the input RGB image, the incidence map and our proposed Camera Image for reference.
3.2 Camera Image Representation
Monocular camera calibration aims to recover the camera intrinsics matrix $\bm{K}$}, which typically composed of four parameters: $f_{x}$, $f_{y}$, $c_{x}$, and $c_{y}$, corresponding to the focal lengths and the optical center coordinates along the $x$-axis and $y$-axis, respectively. This numeric-based representation, however, does not align well with image-based diffusion models, which are primarily designed for generating spatial images. The challenge, therefore, becomes how to effectively leverage powerful pretrained SD models to retrieve implicit camera information.
Figure 3: Error analysis of camera representations. We first use pre-trained VAE to encode and decode each camera representation, and plot the FoV reconstruction errors $(^{\circ})$ here.
To address this challenge, we propose a novel image-based representation, called “Camera Image”, which encodes the camera intrinsic parameters into a 3-channel color image (refer to Fig. 2 for visual representation). This representation preserves most image information, enabling seamless integration into existing diffusion models and requiring only minimal architectural modifications. We reformulate the camera intrinsics into a 2-channel pseudo-spherical representation defined by azimuth $\bm{\theta}$} and elevation $\bm{\varphi}$}. This two-channel formulation enables us to explore the choice of the third channel to prevent mode collapse. Given that VAE encoders typically take a three-channel image as input, it is crucial to determine how to effectively fill the third channel. Simply duplicating one of the existing channels ${\bm{\theta}},{\bm{\varphi}}$ or adding a constant value channel leads to suboptimal results, as detailed in Fig. 3. To enhance the camera representation, we propose a simple yet effective solution by incorporating the grayscale image $\mathbf{g}$} of the input $\mathbf{x}$} into the dense camera representation, reducing the domain gap between the input images and those generated by diffusion models. Consequently, our proposed camera image $\mathbf{c}$} is defined as follows,
${\mathbf{c}}_{(u,v)}=\left[\text{arctan}\left(\frac{r_{1}}{r_{3}}\right),\text{arccos}\left(r_{2}\right),{\mathbf{g}}_{(u,v)}\right],$ (1)
where $\vec{{\bm{r}}}=[r_{1},r_{2},r_{3}]\cong{\bm{K}}^{-1}[u,v,1]^{T}$, $\bm{K}$} is the intrinsic matrix, and $(u,v)$ are the pixel coordinates. $\vec{{\bm{r}}}$ is normalized as a unit vector, and ${\mathbf{g}}_{(u,v)}$ is the gray-scale pixel value sampled at coordinate $(u,v)$. As shown on the right side of Fig. 2, the proposed camera image preserves the high-frequency details of the original scene, making it closely resemble real-world images that diffusion models are designed to process. The incidence map (shown in the middle of Fig. 2) proposed by [23], however, exhibits a large domain gap from the original image domain, which results in suboptimal intrinsics estimations according to our experimental results in Sec. 4.2.
Figure 4: The overview training framework of DM-Calib. The input image $\mathbf{x}$} and the camera image $\mathbf{c}$} are first encoded into latent space using a frozen VAE encoder. We then inject timestamp-relevant noise $\bm{\epsilon}$} into the camera’s latent code, which is concatenated with the image latent code and fed into the subsequent UNet. The UNet is fine-tuned to predict the added noise $\hat{{\bm{\epsilon}}}$.
3.3 Camera Intrinsic Estimation
Diffusion Model for Camera Image Prediction. Our camera intrinsic estimation is built upon the pre-trained latent diffusion model, Stable Diffusion v2.1 [53], which leverages robust image priors trained on the billion-scale LAION-5B dataset [61]. To facilitate the generation of the proposed camera image, we detail our training pipeline in Fig. 4. First, a frozen VAE $\mathcal{E}$ encodes the RGB image $\mathbf{x}$} and its corresponding camera image $\mathbf{c}$} into latent space $\mathbf{z}^{\bm{x}}$ and $\mathbf{z}^{\bm{c}}$, respectively. Multi-resolution noise [32] ${\bm{\epsilon}}^{\bm{c}}$ is then added to the camera latents $\mathbf{z}^{\bm{c}}$, forming the noisy code $\mathbf{z}^{\bm{c}}_{T}$. This code is concatenated with $\mathbf{z}^{\bm{x}}$, serving as the input for the pretrained U-Net. To accommodate our inputs, we double the input channels of the original U-Net and adjust the corresponding parameter weights accordingly. The U-Net is targeted to predict the added noise, and the final loss function is expressed as:
$\mathcal{L}=\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim p_{\text{data}},t\sim p_{t},{\bm{\epsilon}}^{{\bm{c}}}}\left\|\hat{{\bm{\epsilon}}}_{{\bm{\theta}}}(\mathbf{z}^{\bm{c}}_{t};\mathbf{z}^{\bm{x}})-{\mathbf{v}}_{t}\right\|_{2}^{2},$ (2)
where ${\mathbf{v}}_{t}=\alpha_{t}{\bm{\epsilon}}_{t}^{\bm{c}}-\beta_{t}\mathbf{z}^{\bm{c}}_{t}$, and $\bm{\epsilon}_{t}^{\bm{c}}$ is the sampled multiscale noise for the camera image. During inference, we can formulate the generation of the camera image within a generative framework $f\colon{\bm{x}}\in\mathbb{R}^{3}\to\hat{{\bm{c}}}\in\mathbb{R}^{3}$ utilizing v-prediction [56], as follows:
$f(\mathbf{z}^{\bm{x}})=p(\hat{\mathbf{z}}^{\bm{c}}_{T})\prod_{t=1}^{T}p_{\bm{\theta}}\left(\hat{\mathbf{z}}^{\bm{c}}_{t-1}|\hat{\mathbf{z}}^{\bm{c}}_{t}\right),\vskip-5.69054pt$ (3)
where $\mathbf{z}$} is the latent feature and $\hat{\mathbf{z}}^{\bm{c}}_{T}\sim\mathcal{N}({\bm{0}},{\bm{I}})$. After completing the multi-step denoising process using the U-Net, the denoised camera latent representation $\hat{\mathbf{z}}^{\bm{c}}$ is sent to the frozen VAE decoder, yielding the final camera image $\hat{{\mathbf{c}}}={\mathcal{D}}(\hat{\mathbf{z}}^{\bm{c}})$, and we verified that our camera image provides negligible error with respect to the VAE encoder-decoder reconstruction (see Fig. 3). From this generated image, we can extract the numerical representation of the camera intrinsic parameters.
Recover Camera Intrinsics From Camera Image.
With the recoverd camera image $\hat{{\mathbf{c}}}$, camera intrinsic matrix $\bm{K}$} can then be inferred from the first two channels of the camera image via the relation between camera image and camera intrinsic $\bm{K}$} in Eq. 1 :
$\tan(c_{\theta})f_{x}+c_{x}=u,\qquad\frac{1}{\text{cos}(c_{\theta})\text{tan}(c_{\varphi})}f_{y}+c_{y}=v$ (4)
where $\hat{{\mathbf{c}}}_{(u,v)}=[c_{\theta},c_{\varphi},g]$ presents the pixel value of the camera image. Since every two pixels can be used to solve the camera intrinsics, we employ the RANSAC algorithm to align the two lines using all pixel in the camera image. Here, the focal length $f_{x/y}$ and the optical center $c_{x/y}$ are represented as the slope and intercept of the best-fit line of Eq. 4, respectively.
With the proposed camera image and intrinsics estimation, our approach offers applicability to various 3D downstream tasks, including monocular metric depth estimation (MMDE), camera pose estimation, and 3D reconstruction.
3.4 Downstream 3D vision tasks
Monocular Metric Depth Estimation.
To predict metric depth from a single image, the model must possess a deep understanding of the image perspective and estimate accurate intrinsic parameters of the camera. By leveraging the proposed camera calibration method, we repurpose diffusion-based image generators for accurate metric depth estimation. Previous works  [33, 15] mainly investigate affine-invariant depth estimation. However, we find the VAE decoder $\mathcal{D}$ can only predict values in limited range, thus limiting the performance of metric depth estimation. To fix this issue, we formulate stochastic multi-step denoise SD model as one-step deterministic forward process as shown in Fig. 8 of the supplementary. Specifically, we first encode RGB image and our designed camera image $\hat{{\mathbf{c}}}$ via VAE encoder into latent space, noting that no noise is added to both of the latent features. Then, the latent features are sent to the UNet to predict the latent depth features $\hat{{\mathbf{z}}}_{{\bm{d}}}$, and the final depth predictions $\hat{\bm{d}}$ are obtained via the decoder of the VAE. Note that both U-Net $\mathcal{U}$ and the VAE decoder $\mathcal{D}$ are trained to allow predictions in any range. Given the depth labels $\bm{d}$ with its sparse mask $\bm{M}$, the training loss is given by:
$\mathcal{L}_{\text{depth}}=\mathbb{E}_{{\mathbf{x}},\hat{{\mathbf{c}}}\sim p_{\text{data}}}\left\|{\bm{M}}\odot\left[\mathcal{D}(\mathcal{U}(\mathbf{z}^{\bm{x}},\hat{\mathbf{z}}^{\bm{c}}))-{\bm{d}}\right]\right\|$ (5)
This produces clear details and more accurately captures the scene structure, as shown in Sec. 4.3.
Sparse-View 3D Reconstruction &\& Pose Estimation.
Capturing sparse-view images with varying camera settings, particularly focal lengths, complicates object reconstruction using structure-from-motion (SfM) methods like COLMAP [58] due to missing intrinsic parameters and low image overlap. While approaches like DUST3R [73] optimize both intrinsic and extrinsic parameters for reconstruction from sparse viewpoints, they struggle with significantly different intrinsic settings. To address this, we incorporate our estimated intrinsics as a geometry cue for the subsequent reconstruction in Wang et al. [73], fixing the focal length during optimization to demonstrate the robustness of our approach. In Wang et al. [73], the pointmap $X\in\mathbb{R}^{H\times W\times 3}$ is predicted, and the relative pose $P^{*}=\left[R^{*}|t^{*}\right]$ can be recovered via Procrustes alignment [41]. With the povided geometry cue, we achieve better 3D alignment and pose estimation. More disccusion are provided in our appendix.
4 Experiments
[…]
Method Waymo RGBD ScanNet MVS Scenes11 Average
$e_{f}$ $e_{b}$ $e_{f}$ $e_{b}$ $e_{f}$ $e_{b}$ $e_{f}$ $e_{b}$ $e_{f}$ $e_{b}$ $e_{f}$ $e_{b}$
Perspective [30] $\dagger$ 0.444 - 0.166 - 0.189 - 0.185 - 0.211 - 0.239 -
GeoCalib [72] $\dagger$ 0.285 - 0.203 - 0.137 - 0.104 - 0.344 - 0.215 -
WildCame [96] 0.210 0.053 0.097 0.039 0.128 0.041 0.170 0.028 0.170 0.044 0.155 0.041
Unidepth [46] $\ddagger$ - - 0.055 0.052 - - 0.482 $\bm{0.001}$ 0.510 0.051 0.350 0.030
DiffCalib [23] 0.188 0.053 0.092 0.018 0.089 0.041 0.135 0.032 0.108 0.029 0.122 0.030
[…]
Ours $\bm{0.115}$ $\bm{0.036}$ $\bm{0.041}$ $\bm{0.010}$ 0.089 $\bm{0.024}$ $\bm{0.087}$ 0.008 $\bm{0.061}$ $\bm{0.010}$ $\bm{0.078}$ $\bm{0.017}$
4.1 Experimental Setup
Datasets.
For camera intrinsic estimation, the training data is sourced from a variety of datasets, including NuScenes [7], KITTI [19], CityScapes [11], NYUv2 [44], SUN3D [78], ARKitScenes [3], Objectron [1], MVImgNet [86], Hypersim [52], Virtual KITTI [6], Taskonomy [87], and TartanAir [74]. We adopt Waymo [67], RGBD [65], ScanNet [13], MVS [16], and Scenes11 [10] datasets for zero-shot testing.
For metric depth training, we use Taskonomy [87], Hypersim [52], TartanAir [74], Virtual KITTI [6], Waymo [68] and Argoverse2 [76]. Additionally, we incorporate 10k synthetic city samples collected by ourselves. The evaluation is performed on NuScenes [7], ETH3D [59], Diode [70], VOID [77], IBims-1 [35], NYUV2 [44]. For more details, please refer to Sec. 6 of the supplementary.
Evaluation Protocols. For camera intrinsic estimation, we follow the evaluation protocol of [96, 23] using the relative error:
$\begin{split}e_{f}&=\max\left(\frac{|f_{x}^{\prime}-f_{x}|}{f_{x}},\frac{|f_{y}^{\prime}-f_{y}|}{f_{y}}\right)\\ e_{b}&=\max\left(2\cdot\frac{|c_{x}^{\prime}-c_{x}|}{w},2\cdot\frac{|c_{y}^{\prime}-c_{y}|}{h}\right)\end{split}$ (6)
For depth estimation, we use the Absolute mean relative error (A.Rel), the percentage of inlier pixels $\delta_{i}$ with threshold $1.25^{i}$ and scale-invariant error in log scale $\text{SI}_{\text{log}}=100\sqrt{\text{Var }(\varepsilon_{\text{log}})}$.
Baselines. For camera calibration, we compare our method with three non-diffusion based methods, Jin et al. [30], Zhu et al. [96] Piccinelli et al. [46] and Veicht et al. [71], one diffusion-based method [23]. For metric depth estimation, we compare our method with 4 state-of-the-art methods. For additional reference, we also evaluate the generated depth using affine-invariant depth protocols with several affine-invariant depth depth estimation methods.
Implementation Details. Our models are built on the pre-trained Stable Diffusion V2.1 model [53]. To train camera intrinsic estimation model, we employ the AdamW optimizer with a learning rate of $3e^{-5}$ and train the model for 30,000 iterations with a total batch size of 196 on a cluster of 8 Nvidia A800 GPUs. For metric depth estimation, we use the same optimizer and learning rate with a total batch size of 96, and the training process takes approximately 5 days to converge. For all of our downstream 3D vision tasks, we did not use the ground truth camera image but instead relied on intrinsic parameters predicted by our diffusion model.
4.2 Camera Intrinsic Evaluation
We first present our monocular camera calibration results on five zero-shot datasets in Tab. 1. As shown, our method achieves the highest calibration accuracy. Compared to the previous work DiffCalib [23], our approach performs better due to the superior suitability of the proposed Camera Image for diffusion priors, allowing seamless integration with stable diffusion models. Among the methods, Unidepth [46] shows the weakest performance, particularly in real-world, unconstrained scenarios such as the MVS dataset. This subpar result may stem from unbalanced training between the tasks of metric depth estimation and camera intrinsic estimation. Geometry-inspired methods [30, 71] face challenges in achieving strong performance on these datasets as they heavily rely on geometric information, such as vanishing points. Notably, our method also performs well on the highly challenging Scenes11 dataset [10], which features randomly shaped, moving objects, further demonstrating its robustness in extreme conditions.
4.3 Depth Evaluation
Metric Depth Comparison. We evaluate the performance of our method on metric depth estimation task. The quantitative and qualitative results are presented in Tab. 2 and Fig. 5 respectively. Our work significantly outperforms strong baselines such as Metric3D [85] by a large margin, and achieves comparable performance with the SOTA work Unidepth [46]. Based on the visualization results in Fig. 5, compared to UniDepth [46] and Metric3D [85], our method presents sharper details and more accurate structural relationships for the captured scenes.
NuScenes
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ibims
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ETH3D
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RGB GT Ours UniDepth Metric3D $m$
Figure 5: Zero-Shot Metric Depth Estimation Results. We present the predicted metric depth in both various scenes. Our method provides more detailed results and recovers accurate metric.
Table 2: Comparison on Zero-Shot Metric Depth Evaluation. We achieve comparable precision to state-of-the-art models while utilizing less training data.
Method NYU-V2 NuScenes ETH3D DIODE (Indoor) VOID IBims-1
$\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$
iDisc [45] $93.8$ $8.85$ $39.4$ $37.1$ $35.6$ $27.5$ $23.8$ $15.8$ $55.3$ $20.3$ $48.9$ $13.2$
ZoeDepth [4] $90.1$ - $28.3$ $31.5$ $35.0$ $17.6$ $36.9$ $12.8$ $63.4$ $15.9$ $58.0$ $10.9$
Metric3D [85] $92.6$ $9.13$ $72.3$ $29.0$ $45.6$ $18.9$ $39.2$ $11.1$ $65.9$ $16.2$ $79.7$ $10.1$
UniDepth [46] $\bm{97.2}$ $\underline{6.41}$ $\underline{83.3}$ $\underline{22.9}$ $22.9$ $\underline{13.1}$ $\bm{60.4}$ $\bm{9.01}$ $\bm{88.5}$ $\bm{8.26}$ $79.4$ $8.88$
Ours $96.0$ $\bm{6.35}$ $\bm{85.7}$ $\bm{19.8}$ $\bm{49.0}$ $\bm{9.08}$ $\underline{42.2}$ $13.3$ $\underline{84.0}$ $\underline{12.8}$ $\bm{94.4}$ $\bm{7.15}$
Affine-invariant Depth Comparison. Though our method is trained for metric depth, we transform the predicted depth into affine-invariant depth for broader comparisons. As show in Fig. 6, we provide visual results in both in-the-wild and synthetic scenarios. Our approach consistently demonstrates superior spatial structural understanding, such as accurately distinguishing the tree in the background or the Pisa tower, which is correctly inferred to be closer than the nearby church. Quantitative results are presented in Tab. 9 of the supplementary. Despite being designed for metric depth, our model achieves performance comparable to methods tailored for affine-invariant depth.
Input
Marigold
Geowizard
Ours
Figure 6: Zero-shot qualitative affine-invariant depth results. Our method demonstrates superior foreground-background differentiation (e.g., flower) and improved understanding (e.g., wall painting).
4.4 More 3D Vision Tasks
Input
Metric3D
Ours
Figure 7: Metrology of in-the-wild scenes. Our method accurately recovers real-world metrics and demonstrates robustness to variations in focal length.
Monocular 3D Metrology. We evaluate the accuracy of our camera intrinsic estimation and metric depth prediction by estimating the true size of objects captured by cameras with varying focal lengths. For example, using the car shown in Fig. 7, we estimate the distance between the wheels. Compared to Metric3D[85], our method provides more accurate distance estimates across different focal lengths and demonstrates robustness in both outdoor and indoor scenarios (see Fig. 10 of the supplementary). The performance decreases slightly with ultra-wide angles due to the scarcity of small focal length images in our training data, which could be improved by using more balanced datasets.
Sparse-view 3D Reconstruction &\& Pose Estimation. We further demonstrate that our estimated camera intrinsics can be effectively applied to sparse-view 3D reconstruction, especially when photos are taken with varying focal lengths. We evaluate the reconstruction results of Wang et al. [73] on our self-captured images with and without our estimated intrinsics. As shown in Tab. 3, the reconstruction loss, represented by the mean relative distance between corresponding points, was reduced by around 20% on four in-the-wild scenes, confirming the effectiveness of using intrinsic cues for reconstruction. Additionally, as indicated in Tab. 10 of the supplementary, pose estimation performance is also improved when intrinsic cues are used. Quatitative results are shown in Fig. 12 and Fig. 13 of the supplementary, reconstructions without intrinsic cues exhibit notable distortions and misalignments, whereas incorporating intrinsic cues significantly improves accuracy and alignment.
Table 3: Relative distance error. We compare the reconstruction performance with and without intrinsic cues.
Sofa Car Pavilion StoneWall
w/o. cue 1.67 0.87 1.03 1.43
w. cue 1.37 0.68 0.68 1.06
4.5 Ablation Study
Ablation Study on Camera Calibration. We evaluate the effectiveness of our proposed camera image representation and multi-resolution noise strategy through an ablation study on the GSV dataset [2], which includes 20 different camera intrinsic parameters. The study, conducted over 15k iterations, focuses on field-of-view prediction errors. As shown in Tab. LABEL:tab:intrinAbla, the naive approach using $\bm{c}=[\theta,\phi,\theta]$ yields the poorest results due to the domain gap between the generated images and those produced by diffusion models. While multi-resolution noise[32] improves performance slightly, it remains suboptimal. Incorporating our camera image representation significantly reduces the error, and the combination of both strategies produces the best results, demonstrating their complementary effectiveness.
Table 4: Ablation on Intrinsic Estimation.
Multi-Res.
Noise
Camera
Image
[…]
Median Error () $\downarrow$
24.36 25.74
9.10 7.00
6.72 6.33
$\bm{2.54}$ $\bm{2.01}$
Table 5: Ablation on Metric Depth Estimation.
Ablation NYU-v2 KITTI
$\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ A.Rel $\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{SI}_{\log}\downarrow$ A.Rel $\downarrow$
Full Model $\bm{85.8}$ 8.17 $\bm{13.5}$ $\bm{89.1}$ 13.3 $\bm{11.7}$
w.o Real data 26.5 8.80 39.8 69.6 19.0 24.7
w.o One step 77.1 11.9 17.1 76.7 17.1 18.9
w.o Decoder training 76.7 11.1 48.0 83.5 14.5 13.4
w.o Camera Image 83.8 $\bm{8.0}$ 14.6 87.8 $\bm{12.5}$ 12.1
Ablation on Metric Depth Estimation. We investigate the impact of our strategy on metric depth estimation by training a subset of our dataset for 20k iterations, with results shown in Tab. 5. Training solely on synthetic data has proven problematic, especially in indoor scenarios. Our second model, using a traditional multi-step denoising pipeline, integrates both virtual dense and sparse depth data but results in suboptimal performance due to the pipeline’s inability to effectively recognize sparse areas in the ground truth. Additionally, prior methods that froze the VAE decoder during one-step training have shown to be inadequate for metric depth estimation, as demonstrated in our experiments. Omitting the camera image representation slightly reduces accuracy and we further justify the importance of the camera image in metric depth estimation with evaluation on more datasets shown in Sec. 7.8 of the supplementary, demonstrating the indispensable role of the camera intrinsic guidance.
5 Conclusion
In conclusion, we introduced DM-Calib, a diffusion-based framework leveraging our proposed Camera Image for monocular camera calibration. By utilizing the strong priors of stable diffusion models, DM-Calibcan effectively estimate camera intrinsic. Extensive experiments demonstrate its superior performance across a range of 3D vision tasks, consistently outperforming baseline methods in real-world scenarios and varied imaging conditions. Future work could address ultra-wide-angle images by incorporating more diverse training data and improve inference efficiency by developing a few-step diffusion [42] model to further enhance 3D vision tasks.